OpenStudy (anonymous):

find the equation of the tangent to the inverse of the function f(x)=x^2+8x+cos3x at (1,0)

OpenStudy (anonymous):

i need help please.

OpenStudy (anonymous):

Firstly find the equation of the tangent at that point, for the original function. Then find the inverse of that line (I think). I will work it out and check that it is right.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

please try and help me work it out.

OpenStudy (anonymous):

I think I've got it, do you have an answer from the textbook or do you not know the answer?

OpenStudy (anonymous):

i do not know the answer.

OpenStudy (anonymous):

\[y=x^2+8x+\cos(3x)\] \[y'=2x+8-3\sin(3x)\] Now we need to find the derivative at the point, but, since we want the tangent at the point (1,0) on the inverse function, we must inverse this point because we are going to use the derivative of the original function. We do this by swapping the coordinates. (0,1)

OpenStudy (anonymous):

Now, we find the value of the derivative at (0,1) \[y'=2\times0+8-3\sin(3\times0)\] \[y'=8\] Therefore m=8 going through the point (0,1) \[y-y_1=m(x-x_1)\] \[y-1=8(x-0)\] \[y=8x+1\]

OpenStudy (anonymous):

Now the final step is to inverse the tangent. \[y=8x+1\] \[\frac{y-1}{8}=x\] So the tangent of the inverse at (1,0) is: \[y=\frac{1}{8}x-\frac{1}{8}\] Make sense?

OpenStudy (anonymous):

so You don't have to find the inverse of the original function?

OpenStudy (anonymous):

looking through my textbook now i saw the answer to be \[y =\frac{ 1 }{ 8 }\left( x -1 \right)\]

OpenStudy (anonymous):

have understood now . you have got the right answer. I'm grateful

OpenStudy (anonymous):

but what of if you are asked to find the eqn to the tangent of the function f(x)=x^2+8x+sinx at (0,9)

OpenStudy (anonymous):

You take the inverse of the point (0,9); find the equation of the tangent for the original function at that point; inverse the equation for the tangent. It's exactly the same steps, give it a go and if you get stuck then ask for help :)

OpenStudy (anonymous):

the question did not state the inverse of the function but just for the function. i still have to use the inverse?

OpenStudy (anonymous):

Oh, sorry, I didn't see that. You just have to follow the steps without inversing the point and without inversing the tangent equation at the end: Find derivative of function > find equation passing through the specified point

OpenStudy (anonymous):

i have got y=9x+9 i'm i on track

OpenStudy (anonymous):

Are you sure you copied the question correctly? According to this there is no tangent at that point http://www.wolframalpha.com/input/?i=y%3Dx%5E2%2B8x%2Bsinx+tangent+at+%280%2C9%29

OpenStudy (anonymous):

that is what the question stated

OpenStudy (anonymous):

I have to go now. Feel free to use that site as a way to check your answers, and, if you are still stuck then try bumping this question up to the top or making a new question and tagging some of the other helpers online. There is always the chance that the question is written wrong too. Bye-bye

OpenStudy (anonymous):

thank you .

OpenStudy (anonymous):

i wiil use what you taught me to try the inverse of the funtion, if i can get the answer.

OpenStudy (anonymous):

i'm very grateful for your help.

OpenStudy (anonymous):

The inverse will also not work, the only coordinate with a 0 in it that will work for the function you gave me is (0,0). You can try that as a practice and then check it using wolfram if you like. If you liked my help then you can give me a medal by pressing the "best response" button on the right :)

OpenStudy (anonymous):

okay, i have given you the medal

OpenStudy (anonymous):

enjoy your day.